Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Entry  Fri Jun 28 22:49:39 2019, Dirk Hoffmann, Telescope Operations, General, Telescope stuck,  
    Reply  Fri Jun 28 23:06:36 2019, Dirk Hoffmann, Telescope Operations, General, Telescope stuck,  
Message ID: 257     Entry time: Fri Jun 28 22:49:39 2019     Reply to this: 258
Author: Dirk Hoffmann 
Type: Telescope Operations 
Category: General 
Subject: Telescope stuck 
Status:  

After tracking Mk421, we abort current action (acknowledged) and try to go to a new source: RA=308.1, Dec=40.96. Drive Control refuses these parameters ("check your parameters").

General Abort without effect for 2 minutes. Called expert (Ino). He will look into it.

ELOG V3.1.5-fc6679b